### v1

# LLMCheckup: Conversational Examination of Large Language Models via Interpretability Tools

Abstract
Interpretability tools that offer explanations in the form of a dialogue have demonstrated their efficacy in enhancing users’ understanding Slack et al. (2023); Shen et al. (2023), as one-off explanations may occasionally fall short in providing sufficient information to the user. Current solutions for dialogue-based explanations, however, require many dependencies and are not easily transferable to tasks they were not designed for. With LLMCheckup [^1] Code: https://github.com/DFKI-NLP/LLMCheckup Demo: https://youtu.be/ZwN8ZQSXoOU, we present an easily accessible tool that allows users to chat with any state-of-the-art large language model (LLM) about its behavior. We enable LLMs to generate all explanations by themselves and take care of intent recognition without fine-tuning, by connecting them with a broad spectrum of Explainable AI (XAI) tools, e.g. feature attributions, embedding-based similarity, and prompting strategies for counterfactual and rationale generation. LLM (self-) explanations are presented as an interactive dialogue that supports follow-up questions and generates suggestions. LLMCheckup provides tutorials for operations available in the system, catering to individuals with varying levels of expertise in XAI and supports multiple input modalities. We introduce a new parsing strategy called multi-prompt parsing substantially enhancing the parsing accuracy of LLMs. Finally, we showcase the tasks of fact checking and commonsense question answering.
LLMCheckup:
Conversational Examination of Large Language Models
via Interpretability Tools
Qianli Wang1,2   Tatiana $*{}^{*}$1,3   Nils $*{}^{*}$1 Josef van Genabith1,3   Leonhard Hennig1   Sebastian Möller1,2 ${}^{1}$erman Research Center for Artificial Intelligence (DFKI) ${}^{2}$echnische Universität Berlin, Germany ${}^{3}$aarland Informatics Campus, Saarbrücken, Germany {firstname.lastname}@dfki.de
{NoHyper}*Equal contribution
Figure 1: LLMCheckup dialogue with data augmentation and rationalization operations on a commonsense question answering task (ECQA). Boxes (not part of the actual UI) indicate the original instance from the dataset as well as its prediction (cyan) and the explanation requested by the user (orange).
1 Introduction
To unravel the black box nature of deep learning models for natural language processing, a diverse range of explainability methods have been developed Ribeiro et al. (2016); Madsen et al. (2022); Wiegreffe et al. (2022). Nevertheless, practitioners often face difficulties in effectively utilizing explainability methods, as they may not be aware of which techniques are available or how to interpret results provided. There has been a consensus within the research community that moving beyond one-off explanations and embracing conversations to provide explanations is more effective for model understanding Lakkaraju et al. (2022); Feldhus et al. (2023); Zhang et al. (2023) and helps mitigate the limitations associated with the effective usage of explainability methods to some extent Ferreira and Monteiro (2020); Slack et al. (2023).
Figure 2: On the left, a dialogue example asking for explanation in natural language about a ECQA-like customized question. The workflow of LLMCheckup is shown on the right side.
In the field of NLP, two dialogue-based interpretability tools, InterroLang Feldhus et al. (2023) and ConvXAI Shen et al. (2023), have been introduced. Both tools employ multiple, separately fine-tuned LMs to parse user intents and dedicated external LMs to provide explanations.
By contrast, our system, LLMCheckup, only requires a single LLM and puts it on “quadruple duty”: (1) Analyzing users’ (explanation) requests (§2.1, §5.2), (2) performing downstream tasks (§4), (3) providing explanations for its outputs (§3), and (4) responding to the users in natural language (§2.3). Instead of using many LMs to explain the behavior of another LLM, LLMCheckup allows us to directly request the same LLM used for parsing to self-explain its own behavior. Any LLM can be deployed “out of the box” and confronted with diverse NLP tasks, explanation generation and intent recognition. LLMCheckup provides a unified and compact framework, effectively consolidating and obviating the requirements for additional LMs to generate explanations.
2 LLMCheckup
LLMCheckup is an interface for chatting with any LLM about its behavior. We connect several interpretability tools (§3), s.t. self-explanations take model internals, datasets and documentation into account. User requests for explanations are recognized via a text-to-SQL task performed by the LLM (§2.1).
2.1 System architecture
Figure 2 illustrates the interaction flow of LLMCheckup. When a user asks a question on the user interface, the question will be parsed as an SQL-like query by the LLM in use. E.g., the first user question in Figure 1 would be parsed as "filter id 26 and rationalize". The corresponding operation (see Table 6) will then be matched (parsing) and executed. Finally, the generated explanation will be filled into a template that will be returned back to the user Slack et al. (2023); Feldhus et al. (2023) (response generation).
Parsing
To enhance the recognition of users’ intentions, we employ two methods: Guided Decoding and Multi-prompt Parsing.
Guided Decoding (GD) ensures that the generated output adheres to predefined grammatical rules and constraints Shin et al. (2021) and that generated parses align with predefined operation sets Slack et al. (2023). GD is generally more suitable for smaller LMs.
As an alternative to GD, we also propose and implement a novel Multi-prompt Parsing (MP) approach. While GD pre-selects prompts based on the embedding similarity with user input, the model does not see all the available operations at once and the pre-selection may not include the examples for the actual operation required. With MP, we test whether showing all possible operations in a simplified format as the first step (i.e., without any attributes such as the instance ID or the number of samples) and then additionally prompting the model to fill in more fine-grained attributes can improve performance. We provide further details about MP in Appendix B. §5.2 evaluates the performance of both parsing approaches.
2.2 Key features
Supported NLP models
We identify five auto-regressive LLMs representative of the current state-of-the-art in open-source NLP (as indicated in the left column of Table 2) available through Hugging Face Transformers Wolf et al. (2020). The diverse choice of models demonstrates that our system is generalizable and that in principle it is theoretically possible to choose any Transformer-type model to be explained. While Falcon-1B Penedo et al. (2023) and Pythia-2.8B Biderman et al. (2023) are available for users with limited hardware resources (RAM/GPU), it is generally not recommended to use them due to their small model size, which may negatively affect the user’s perception of the performance. Llama2-7B Touvron et al. (2023) and Mistral-7B Jiang et al. (2023) are both mid-sized with 7B parameters, while Stable Beluga 2 Mukherjee et al. (2023) is a fine-tuned version of Llama2-70B. To facilitate the deployment of large models in a local environment, LLMCheckup offers support for various forms of LLMs. This includes the usage of quantized models through GPTQ Frantar et al. (2023), loading models in 4-bits with the assistance of bitsandbytes Dettmers et al. (2022), and the implementation of a peer-to-peer solution using Petals Borzunov et al. (2023), enabling efficient deployment on a custom-level GPU.
Tutorial
To help laypeople get background knowledge in XAI, we introduce a tutorial functionality. It is based on prompting with different roles corresponding to levels of expertise in XAI (Figure 3) and enables us to provide tailored meta-explanations of supported operations to individuals. For example, for the beginner level, we add a system prompt hinting at the expertise: “As a NLP beginner, could you explain me the question: What is data augmentation?” (Figure 4).
External information retrieval
LLMs may sometimes generate responses that appear plausible but are actually incorrect predictions known as hallucinations Welleck et al. (2020). LLMCheckup provides users with the capability to access information by conducting search through external knowledge bases, promoted by the integration of googlesearch [^2] https://github.com/Nv7-GitHub/googlesearch (Figure 5). Users can cross-reference the retrieved information with the provided explanations, thereby achieving a more comprehensive understanding.
Customized inputs & prompts
In comparison to TalkToModel Slack et al. (2023), which was limited to three datasets, LLMCheckup offers users the freedom to enter custom inputs according to their preferences, going beyond just querying instances from the provided dataset. In addition, inspired by PromptSource Bach et al. (2022), users can insert pre-defined prompts to the explanation step (§3.2), but the Prompt Editor also allows fully customized prompts, allowing for a personalized and tailored experience. All custom inputs will be saved and can be inspected and reused later on via a dedicated custom input history viewer.
Multi-modal input format
Motivated by Malandri et al. (2023), LLMCheckup not only accepts text input from users but also provides support for other modalities such as images and audio. To facilitate this, we employ specific packages and models tailored to each modality. For optical character recognition (OCR), we use EasyOCR [^3] https://github.com/JaidedAI/EasyOCR. For audio recognition, we employ a lightweight fairseq S2T [^4] https://huggingface.co/facebook/s2t-small-librispeech-asr model Wang et al. (2020) trained on Automatic Speech Recognition (ASR).
Suggestion of follow-up questions
To guide the user through the conversation, we implemented a suggestions mode. When activated, the user will get suggestions for related operations that LLMCheckup can perform (see Appendix C for more detail), e.g., if the user asks about the top $k$ attributed tokens for a specific sample, they will get a suggestion to have a look at the rationales generated by the model since both token attributions and rationalizations belong to the “Explanation" category also displayed in the user interface. Additionally, for each generated suggestion we check whether it already appears in the dialogue history to make sure that the user does not receive repetitive suggestions for the operations that they have already been performed. E.g., if the user inquires about the counterfactual operation and the model explains how it works, LLMCheckup will store this information and will not suggest explaining counterfactuals again.
Dialogue sharing
LLMCheckup offers the functionality to export the dialogue history between the user and the deployed LLM as a JSON file that contains the user’s questions and the corresponding generated responses. This simplifies data collection and sharing of conversation logs between users.
2.3 Interface
LLMCheckup provides a user interface including a chat window to enter questions and settings and is implemented in Flask and can be run as a Docker container. LLMCheckup builds on top of the chat window from TalkToModel Slack et al. (2023) for continuous interaction as well as buttons for question generation and the dataset viewer from InterroLang Feldhus et al. (2023), which facilitates dataset exploration and generates specific questions for all available operations to inspire users to come up with their own questions. Figure 3 shows the LLMCheckup interface.
In comparison to TalkToModel Slack et al. (2023) and InterroLang Feldhus et al. (2023), our interface provides the following new components:
  • Checkbox for selecting XAI user expertise level;
  • Input field for custom inputs;
  • Prompt editor including a checkbox for selecting additional prompts and a display of the currently used system prompt;
  • Support for multimodal input through buttons for image upload and audio recording.
Dataset COVID-Fact ECQA
Model Size Consistency Claim Fluency Evidence Fluency Consistency Question Fluency
Mistral 7B 0.66 0.88 0.96 0.50 0.76
Llama2 7B 0.65 0.88 0.94 0.50 0.76
Stable Beluga 2 70B 1.00 0.85 0.96 1.00 0.73
Table 1: Consistency and fluency scores of data augmentation from three models. falcon and pythia are not considered due to poor performance because of small model size.
3 NLP explainability tools
Table 3 and Table 4 show examples of explanations for each supported explainability method by LLMCheckup.
3.1 White-box
We provide white-box explainability operations that are gradient-based or embedding-related, both of which are not easily applicable to the prompting paradigm of LLMs.
Feature attribution
Feature attribution methods quantify the contribution of each input token towards the final outcome. In LLMCheckup, we deploy various auto-regressive models (§2.2), for which Inseq Sarti et al. (2023) is used to determine attribution scores. We support representative methods from Inseq, including Input x Gradient Simonyan et al. (2014), Attention Bahdanau et al. (2015), LIME Ribeiro et al. (2016), and Integrated Gradients Sundararajan et al. (2017) [^5] Details on the Inseq integration are described in App. E..
Embedding analysis
By calculating the cosine similarity between the sentence embeddings of the instances in datasets, we can retrieve relevant examples Cer et al. (2017); Reimers and Gurevych (2019) and present them for contextualizing the model behavior on the input in question.
3.2 Black-box
Data augmentation
Augmentation involves synthesizing new instances by replacing text spans of the input while preserving the semantic meaning and predicted outcomes Ross et al. (2022). As a baseline, NLPAug is used to substitute input by WordNet’s Miller (1995) synonyms [^6] https://github.com/makcedward/nlpaug NLPAug offers various techniques for generating augmented text, including back-translation, word embedding substitution, etc.. However, we constrain the augmentation to only utilize WordNet to avoid additional LMs.. Alternatively, data augmentation can be achieved by providing the LLM with a few demonstrations as prompts Dai et al. (2023).
Counterfactual generation
Unlike data augmentation, counterfactuals are about input edits causing the predicted outcome to be different Wu et al. (2021); Chen et al. (2023). Counterfactuals are generated by prompting LLMs with manually crafted demonstrations.
Rationalization
Rationalization aims to provide free-text explanations that elucidate the prediction made by the model Camburu et al. (2018); Wiegreffe et al. (2022) (an example is shown in Figure 1). The use of Chain-of-Thought (CoT) prompting enhances the reasoning capabilities of LLMs by encouraging the generation of intermediate reasoning steps that lead to a final answer Wei et al. (2022); Wang et al. (2023b). Different CoT strategies can be applied depending on users’ preferences, including Zero-CoT Kojima et al. (2022), Plan-and-Solve Wang et al. (2023a), and Optimization by PROmpting (OPRO) Yang et al. (2023) (Figure 3).
4 Use cases
In this paper, we demonstrate the usefulness of LLMCheckup with two typical NLP tasks: Fact checking and commonsense question answering.
4.1 Fact checking
The importance of fact checking has grown significantly due to the rapid dissemination of both accurate information and misinformation within the modern media ecosystem Guo et al. (2022). COVID-Fact Saakyan et al. (2021) is a dataset resembling FEVER Thorne et al. (2018) that encompasses various claims, supporting evidence for those claims, and contradictory claims that have been debunked by the presented evidence.
4.2 Commonsense question answering
Unlike question answering, commonsense question answering (CQA) involves the utilization of background knowledge that may not be explicitly provided in the given context Ostermann et al. (2018). The challenge lies in effectively integrating a system’s comprehension of commonsense knowledge and leveraging it to provide accurate responses to questions. ECQA Aggarwal et al. (2021) is a dataset designed for CQA. Each instance in ECQA consists of a question, multiple answer choices, and a range of explanations. Positive explanations aim to provide support for the correct choice, while negative explanations serve to refute incorrect choices. Additionally, free-text explanations in natural language are included as general justifications.
5 Evaluation
5.1 Data augmentation
We assess the quality of the generated augmented output based on two key aspects: (1) consistency: the metric represents the proportion of instances where the augmentation process does not lead to a change in the label Li et al. (2023a); Dai et al. (2023); (2) fluency: assesses how well the augmented output aligns with the original data in terms of semantic similarity Ross et al. (2021) measured by SBERT [^7] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Table 1 indicates that Mistral and Llama2 exhibit comparable performance, while Stable Beluga 2 displays substantially higher consistency scores on two tasks, although it may exhibit lower fluency in certain cases. The overall performance on ECQA is relatively lower compared to COVID-Fact. This difference in performance can be attributed to the increased complexity and difficulty of the ECQA task.
5.2 Parsing
To assess the ability of interpreting user intentions by LLMs, we quantify the performance of each deployed model by calculating the exact match parsing accuracy Talmor et al. (2017); Yu et al. (2018) on a manually created test set, which consists of a total of 119 pairs of user questions and corresponding SQL-like queries. As an additional baseline, we employ the nearest neighbor approach that relies on comparing semantic similarity.
We assess parsing accuracy of our two approaches, GD and MP (§2.1). Table 2 shows that, as model size increases, the parsing accuracy tends to increase and MP demonstrates a notable improvement over GD. Despite Stable Beluga 2 having a larger size compared to 7B models, its parsing performance only marginally surpasses that of Mistral and Llama2. This can be partially attributed to the difficulty of the parsing task [^8] We have a total of 21 operations displayed in Table 6, and many of these offer multiple options. For instance, score operation supports $F_{1}$, precision, recall and accuracy matrices. and the number of demonstrations, as larger models may require a greater number of demonstrations to fully comprehend the context Li et al. (2023b). Parsing results are discussed in more detail in Appendix B.
Model Size Strategy Accuracy
Nearest Neighbor - - 42.24
Falcon 1B GD 47.41
Pythia 2.8B GD 51.72
Llama2 7B GD 64.71
Mistral 7B GD 55.88
Stable Beluga 2 70B GD 67.23
Falcon 1B MP 64.15
Pythia 2.8B MP 75.91
Llama2 7B MP 82.35
Mistral 7B MP 84.87
Stable Beluga 2 70B MP 88.24
Table 2: Exact match parsing accuracy (in %) for different models. GD = Guided Decoding prompted by 20-shots; MP = Multi-Prompt parsing.
6 Discussion
We highlight the broad generalizability of LLMCheckup concerning diverse NLP tasks and models. Our approach empowers any auto-regressive LLM to handle dialogue-based explainability including parsing and explanation generation. Aside from the NLP tasks showcased in Section 4, LLMCheckup can be easily expanded to accommodate other tasks not directly related to classification, such as summarization or translation [^9] For other NLP tasks, the predict operation should be replaced by a task-specific operation and all explanation operations can be retained and easily adapted.. In contrast to ConvXAI Shen et al. (2023) and InterroLang Feldhus et al. (2023), which require a fine-tuned model for each specific use case, LLMs used in LLMCheckup possess remarkable zero-/few-shot capabilities Brown et al. (2020), which allows them to effectively handle many tasks without requiring fine-tuning. Although the quality of an explanation could be enhanced with further fine-tuning, LLMCheckup uses model outputs out of the box.
7 Related work
Interfaces for interactive explanations
LIT Tenney et al. (2020) is a GUI-based tool available for analyzing model behaviors across entire datasets. However, LIT has less functionalities in terms of prompting and lower accessibility, e.g. no tutorial and a lower level of integration with HuggingFace. CrossCheck Arendt et al. (2021) exhibits the capability to facilitate quick cross-model comparison and error analysis across various data types, but adapting it for other use cases may need a substantial amount of code modification and customization. XMD’s Lee et al. (2023) primary purpose is model debugging, but it shares similarities in the focus on feature attributions, visualization of single instances and user feedback options. It is, however, limited to feature attribution explanations and smaller, efficiently retrainable models. IFAN Mosca et al. (2023) enables real-time explanation-based interaction with NLP models, but is limited to the sequence-to-class format, restricting its applicability to other tasks. Furthermore, IFAN offers a limited set of explainability methods, consisting of only one local and one global method.
Dialogue-based systems for interpretability
Carneiro et al. (2021) point out that conversational interfaces have the potential to greatly enhance the transparency and the level of trust that human decision-makers place in them. According to Zhang et al.’s user studies, delivering explanations in a conversational manner can improve users’ understanding, satisfaction, and acceptance. Jacovi et al. (2023) emphasizes the necessity of interactive interrogation in order to build understandable explanation narratives. ConvXAI Shen et al. (2023), TalkToModel Slack et al. (2023), InterroLang Feldhus et al. (2023) and Brachman et al. (2023) share some similarities with our system, but are more complex in their setup and consider fewer explainability methods. Additionally, they might overrely on external LMs to explain the deployed model’s behavior, whereas LLMCheckup places a strong emphasis on self-explanation, which is crucial for faithfulness. Finally, LLMCheckup uses auto-regressive models, as they have become increasingly dominant in various NLP applications nowadays. In iSee Wijekoon et al. (2023), a chatbot adapts explanations to the user’s persona. Explanation strategies and user experiences can be captured and shared. The system is currently still in development, but shares many functionalities, except that LLMs are not in their scope yet.
8 Conclusion
We present the interpretability tool LLMCheckup, designed as a dialogue-based system. LLMCheckup can provide explanations in a conversation with the user facilitated by any auto-regressive LLM. By consolidating parsing, downstream task prediction, explanation generation and response generation within a unified framework, LLMCheckup streamlines the interpretability process and serves as a baseline for future investigation.
Future work includes exploring RAG models Lewis et al. (2020) combined with explainability, as currently LLMCheckup relies on search engines for external information retrieval. We also want to add multi-modal models, so converting image or audio input to texts would no longer be necessary, but the current state of interpretability on such models lags behind unimodal approaches Liang et al. (2023). Integrating our system into HuggingChat [^10] https://huggingface.co/chat/ would further increase the visibility and accessibility through the web.
Limitations
In LLMCheckup, we do not focus on dataset analysis or data-centric interpretability, but on how a model responds to single inputs. There are a lot of practical cases, e.g. medical report generation Messina et al. (2022), gender-aware translation Attanasio et al. (2023), where users are not interested in raw performance metrics on standard benchmarks, but are interested in detecting edge cases and investigating a model’s behavior on custom inputs.
English is the main language of the current system. Multilinguality is not supported, as both the interface, the responses, tutorial and the explained models are monolingual. While it would be possible to adapt it to other languages by translating interface texts and prompts and using a model trained on data in another target language or multiple ones, it remains to be seen to which extent multilingual LLMs can do quadruple duty as well as the current model does for English.
In LLMCheckup, users have the flexibility to input data in different modalities, including images and audio. However, for audio and images, LLMCheckup will convert the audio content and texts contained within the images into textual format for further processing and analysis. Besides, the explanations and responses generated by our system are currently limited to the text format – apart from the heatmap visualization of feature attribution explanations.
The QA tutorial only aims to provide explanations for supported operations in XAI to individuals with different levels of expertise. However, the explanations, e.g. rationales, generated by the LLM may not inherently adapt to users’ specific expertise levels Zhang et al. (2023). In the future, we will explore how to prompt the models to provide simple explanations reliably.
In LLMCheckup, we employ a single LLM to serve quadruple-duty simultaneously. However, models with lower parameter counts may exhibit limitations in certain types of explanation generation, particularly when using prompting techniques like rationalization or counterfactual generation Marasovic et al. (2022).
Appendix A Explanation examples
User LLMCheckup
Hi! What do you predict for instance with id 2435?
10 most important features for it regarding attention?
And starting it how would a new instance look like?
Can you explain it in natural language?
Table 3: Sample dialogues for welcome words, prediction (predict), feature attribution (nlpattribute), data augmentation (augment) and rationalization (rationalize) for the ECQA use case.
User LLMCheckup
How would you flip the prediction for it?
Is it possible to retrieve an example that is similar to it?
Table 4: Sample dialogues for counterfactual (nlpcfe), similar (similar) for the ECQA use case.
Table 3 and Table 4 displays examples of explanations for each supported explainability method. In each screenshot, the operation name is highlighted in blue.
Appendix B Multi-prompt parsing
As a first step, MP queries the model about the main operation (e.g., feature attribution, counterfactual generation, dataset statistics etc., see the full list of the available operations in Appendix D). Next, depending on the chosen operation, MP selects the operation-specific prompts with 2-7 demonstrations [^11] The number of demonstrations depends on the difficulty of operation, e.g., how many attributes it may have. (user query and parsed outputs examples) to generate the full parses that may include several attributes. E.g., for the user input "What are the feature attributions for id 42 based on the integrated gradients?" we start by generating nlpattribute and then augment the parse with the second prompt and transform it into filter id 42 and nlpattribute integrated_gradient.
Since the output of the model is not constrained compared to GD, in the MP setting, we need to check whether the model’s output matches any of the available operations and if there is no exact match we employ SBERT and compare the embedding similarity between the user input and the operations of LLMCheckup (e.g., sometimes the model outputs synonyms but we need to match the operation names exactly to ensure the correct parsing). Additionally, we implement some checks to avoid possible hallucinations, e.g., if the model outputs an id that is not present in the user input we remove it from the parsed text.
Model #max_new_tokens Accuracy
Falcon 10 64.15
Falcon 20 64.15
Pythia 10 75.91
Pythia 20 63.03
Llama2 10 74.79
Llama2 20 82.35
Llama2-GPTQ 10 82.63
Llama2-GPTQ 20 87.40
Mistral 10 84.87
Mistral 20 71.43
Mistral-GPTQ 10 78.71
Mistral-GPTQ 20 68.91
Stable Beluga 2 10 88.24
Stable Beluga 2 20 86.55
Table 5: Parsing accuracy (in %) using MP with different number of maximum new tokens. Note that for the Llama2-7b and Mistral-7b models, we offer various options for quantization. In this case, we have chosen GPTQ as the representative method.
Table 5 summarizes our parsing evaluation results for different models with different number of ’max_new_tokens‘ for generation specified in the configuration file. We found that this parameter is very important for the performance of the MP parser. Pythia-2.8B, Mistral-7B and Stable Beluga 2 achieve better scores with less tokens to generate (10) and Llama2-7B shows better performance with larger number of tokens (20). The only model that is unaffected by the ’max_new_tokens‘ parameter is Falcon-1B which also has the smallest size. The majority of the models perform better when allowed to generate only brief texts and if they are given more tokens to generate they tend to hallucinate and detract from the actual parsing task.
Originally, we treated all models in the same way during parsing, however, after looking at some generated outputs we realized that Falcon-1B and Pythia-2.8B are not good at extracting ids and often can only recognize the main operation. Hence, for these models we have an additional step that extracts a potential id from the user input and adds it to the parsed operation.
As expected, larger models tend to perform better than the ones with less parameters. However, we also found that the quantized Llama model outperforms its full (non-quantized) version on the parsing task.
Appendix C Follow-up suggestions
We group the suggestions into several categories that include metadata operations (e.g., dataset statistics, model types etc.), prediction-related operations (e.g., predict, count or show mistakes), explanation-based operations (e.g., attributions for top $k$, attention scores and integrated gradients or free-text rationale), NLU (similarity and keywords) and input perturbations (counterfactuals and data augmentation).
The user always has an option to decline the suggestion and ask something different. We check whether the user agrees to the LLMCheckup suggestions by computing the similarity scores between the input and the confirm/disconfirm templates with SBERT.
Appendix D Supported operations
We define an operation as a mapping between (a part of) an SQL-type query and a function executed by LLMCheckup. Table 6 enumerates all operations facilitated by LLMCheckup, including all explainability methods mentioned in §3 and other supplementary operations.
Filt. filter(id) Access single instance by its ID
includes(token) Filter instances by token occurrence
Prediction $*{}^{*}$ Get the prediction of the given instance
randompredict(number) Precompute a subset of instances at random
mistakes(subset) Count number of wrongly predicted instances
score(subset, metric) Determine the relation between predictions and labels
Data show(list) Showcase a list of instance
countdata(list) Count number of instances within the given list
label(dataset) Describe the label distribution across the dataset
Meta data() Information related to the dataset
model() Metadata of the model
About function() Inform the functionality of the system
self() Self-introduction
qatutorial(op_name) Provide explanation for the given operation
Logic and(op1, op2) Concatenation of multiple operations
or(op1, op2) Selection of multiple filters
Explanation nlpattribute(inst., topk, $*{}^{*}$ Provide feature attribution scores
$*{}^{*}$ Explain the output in natural language
$*{}^{*}$ Generate counterfactuals
similarity(instance, $*{}^{*}$ Output top $k$ similar instances in the dataset
$*{}^{*}$ Augment the input text
Table 6: Operations marked with ($*{}^{*}$) support the use of custom inputs.
Appendix E Details on feature attribution
In LLMCheckup, we do not exhaustively employ all Inseq’s methods for feature attribution. Instead, we selectively choose certain representative methods from our perspective. Nevertheless, we would like to emphasize that it is straightforward to incorporate addition methods such as Saliency Simonyan et al. (2014), Occlusion Zeiler and Fergus (2014), Sequential Integrated Gradients Enguehard (2023).
Appendix F Interface
Figure 3 display the user interface of LLMCheckup. Within the chat window, a user’s question requesting the explanation for instance with ID 4180 in natural language is shown, followed by a response that includes the corresponding rationale. On the right side of the window, there is a prompt modification section where the Optimization by Prompting (OPRO) strategy is chosen for rationalization. The icons associated with each strategy describe them in detail when they are hovered over, including the corresponding prompts that can be appended after the system prompt we have designed.
Figure 3: LLMCheckup interface with initial welcome message, generated free-text rationale and sample generator buttons. Expert level of knowledge in XAI is selected and Optimization by Prompting (OPRO) strategy is used. For an example dialogue, see Table 3 and Table 4.
Appendix G QA tutorial
Figure 4 shows tutorials for data augmentation with different level of expertise in XAI.
(a) QA Tutorial for data augmentation with beginner level of knowledge in XAI.
(b) QA Tutorial for data augmentation with expertise level of knowledge in XAI.
(c) QA Tutorial for data augmentation with expert level of knowledge in XAI.
Figure 4: QA tutorial with different knowledge level in XAI.
Appendix H External information retrieval
Figure 5 shows the external information retrieval for an instance from COVID-Fact.
Figure 5: External information retrieval of an instance from COVID-Fact.